Suppose John’s demand for a good is given by P=200-1/5 Q and that marginal revenue is MR=200-0.4Q. Marginal cost and average total cost are constant at $20 per unit.
Consider the following quantity discount.
The price for the first 400 units a customer purchases is $120 per unit, but any additional units have a price of $80 per unit. How many units will John purchase?
1
Expert's answer
2020-06-11T10:54:16-0400

John will purchase where the marginal revenue is equal to the marginal cost.


Equating the marginal revenue to the marginal cost:



"200 - 0.4Q = 20 \\[0.3cm]
0.4Q = 180\\[0.3cm]
Q = 450"

Without the quantity discount, John will purchase 450 units.


The price without quantity discount is:



"P = 200 - \dfrac{1}{5}(450)\\[0.3cm]
P = \$110"

Without the quantity discount, John will spend:



"\$110\times 450=\$49,500"

The quantity discount is that John gets to pay $120 for the first 400 units. Thus, for 450 units, his spending will be:


"(\$120\times 400) + \$80(450 - 400) = \$52,000"


Therefore, John will buy just 450 units.